# Energy-Efficient Approximate Full Adders Applying Memristive Serial IMPLY Logic For Image Processingthanks: Preprint Submitted to arXiv

### v1

@@ 4.2 Error analysis simulation and comparison
Two items should always be evaluated and analyzed in approximate computing. The first item, the circuit analysis metrics improvement, is analyzed in the last subsection. The computation’s accuracy, as the second main item in approximate computing, is analyzed in this subsection. Estimating the error magnitude of an error-tolerant application’s outputs is possible by examining error analysis metrics [42]. Error analysis metrics such as MED and NMED can be applied to analyze the proposed circuits’ accuracy. The smaller the error analysis metrics, the higher the computations’ accuracy. All possible input combinations (65536 different input patterns) are applied to the 8-bit approximate adders of scenarios 1-3 designed based on ICIS1-3, ECIS, SIAFA1-4 [3], and SAFAN [31] to calculate the error analysis metrics. The results of error analysis metrics are shown in Table 15.
@@
The main goal of proposing the ECIS and overcoming its hardware complexity compared to the ICIS1-3 is to prevent the spread of inexact $C_{out}$ from LSBs to the MSBs and increase the accuracy of computations. Error analysis metrics (e.g., MED and NMED) are minimum when the ECIS is applied in the structure of scenarios 1-3, according to the results of Table 15. The accuracy of computations increases when the ECIS full adder is applied in larger arithmetic structures compared to the ICIS1-3, SIAFA1-4 [3], and SAFAN [31] based on the results of Table 15. So, the primary purpose of proposing the ECIS is achieved. The error analysis metrics of ICIS2 and ICIS3 full adders are equal, based on their truth tables, in scenarios 1-3, and the results are reported in Table 15. The ED of these two approximate full adders equals 3, and their computations’ accuracy is lower than ECIS (ED=2). The computations’ accuracy decreased by 9%, 2%, and 1%, respectively, by applying the ICIS2 and ICIS3 instead of SIAFA1 [3] and SIAFA3 [3] in the structures of scenarios 1-3. It should be mentioned that the number of computational cycles has also decreased by 5%, 7%, and 9% by applying the ICIS2 and ICIS3 instead of SIAFA1 [3] and SIAFA3 [3] in these three scenarios, respectively. The computations’ accuracy of ICIS1 is lower than the other proposed approximate full adders in this article, SIAFA1 [3], and SIAFA3 [3] in scenarios 1-3. The accuracy of computations decreased from 4%-10% by applying ICIS1 instead of SIAFA1 [3] and SIAFA3 [3] in scenarios 1-3, but the computational steps improved by 5%-9% in these scenarios. This approximate full adder increases the accuracy of computations in scenarios 1-3 compared to SIAFA2 [3], SIAFA4 [3], and SAFAN [31].
@@ 4.3 Application-level simulation and comparison
The reduction of circuit complexity in approximate computing should be accompanied by acceptable accuracy reduction so that this computational method can be effectively applied in error-resilient applications. Image processing as an error-resilient application is a complex and data-intensive application widely applied in daily human lives [3]. This application is a suitable candidate for in-memory approximate computing by applying memristors [3, 9]. So, the proposed approximate full adders are applied in the computational structures of scenarios 1-3 and are simulated behaviorally in three different image processing applications of image addition, motion detection, and grayscale filter. PSNR, SSIM, and MSSIM are applied to evaluate the output image quality metrics. The proposed full adders have equal single-bit error analysis metrics but do not have equal output image quality metrics in image processing applications. The reason is that the input distributions in images differ from those in the error analysis simulations [3].
@@ 4.3.1 Image addition application
According to Table 16, the highest quality of the output images is obtained when ECIS is applied in the computational structures of scenarios 1-3 introduced in subsection 4.1. The lowest quality of the output images (PSNR) corresponded to SIAFA2 [3] and SAFAN [31]. Also, the quality of output images is acceptable when ICIS1-3 are applied in the computational structures of scenarios 1-3 (PSNR is higher than 30 dB). The output image quality is higher than 45 dB when the ECIS is applied in scenario 1. The quality of the output images varied between 43.94 dB and 44.5 dB when ICIS1-3, SIAFA1 [3], and SIAFA3 [3] are applied in the computational structure of scenario 1. The quality of the output images is higher than 39 dB only when ECIS is applied in the second scenario’s computing structure. The quality of the output images is between 38-39 dB when ICIS1-3, SIAFA1 [3], and SIAFA3 [3] are applied in the second scenario’s RCA structure. The outputs of the image addition application have the least image quality when SAFAN [31] and SIAFA2 [3] are applied in the second scenario’s arithmetic architecture (PSNR is less than 38 dB). The quality of output images is higher than other circuits by applying ECIS and ICIS3 in the third scenario (PSNR is higher than 33 dB). The outputs are acceptable (PSNR is higher than 30 dB) when ICIS1, ICIS2, SIAFA1 [3], SIAFA3 [3], SIAFA4 [3], and SAFAN [31] are applied in the LSBs of the third scenario’s approximate adder. Suppose the number of approximate full adders (ECIS, ICIS1-3, SIAFA1-4 [3], and SAFAN [31]) increases in the LSBs of an 8-bit RCA to six bits. In that case, the image quality of the outputs generated by these RCAs is less than 30 dB, and the output images are unacceptable.
@@ 4.3.2 Motion Detection Application
Two consecutive frames are placed in the subfigures 12a and 12b. The result of the subtraction of these two consecutive frames applying an exact 8-bit adder is shown in the subfigure 12c. Subfigures 12d-12g are the output images obtained from the subtraction of these two consecutive frames applying the proposed approximate full adders in the RCA of scenario 2. The image quality metrics of the output images in different scenarios are written in Table 17.
@@
According to the results of Table 17, ECIS and ICIS1-3 are among the first six approximate full adders in the application of motion detection in all three scenarios. When the proposed circuits are applied in the computational structures of scenarios 1-3, the quality of the output images is higher than 45 dB, 40 dB, and 34.9 dB, respectively. The image quality metrics of the outputs of motion detection application are not acceptable if the number of approximate full adders in the LSBs of an 8-bit approximate adder structure is changed from 5 to 6 (only the two MSBs are computed applying the exact full adder). The PSNR of output images is higher than 29,5 dB, but the motions could not be detected by applying all four proposed approximate full adders in this structure. The same results are obtained by applying the SOA [3, 31] in this RCA, and it is impossible to detect motion by applying these approximate full adders.
@@ 4.4 Analysis of the trade-off between Circuit and Error evaluation metrics
The proposed circuits were analyzed only by circuit analysis metrics in subsection 4.1. Different error analysis criteria and image quality metrics were applied to analyze the proposed circuits in subsections 4.2, and 4.3. The reduction of accuracy in computations leads to the reduction of circuit complexity in approximate computing. We compared the proposed circuits by simultaneously considering the circuit and error evaluation criteria in this subsection. The $FOM$ introduced in this article considers the $\underline{E}$nergy consumption and computational $\underline{D}$elay $\underline{P}$roduct (EDP) as circuit analysis metrics and the accuracy-related metrics of 1-NMED and PSNR according to the third scenario’s results.
This $FOM$ is defined as
$FOM=\frac{EDP}{(1-NMED)\times PSNR}$ (27)
The results of $FOM$ for the proposed circuits, SIAFA1-4 [3], and SAFAN [31] are written in Table 19. The average PSNR value of three 8-bit image processing applications assessed in the last subsection is considered to calculate the $FOM$. ICIS3, ICIS2, and ICIS1 made the best trade-off between the circuit complexity and the accuracy of the computations, respectively, according to Table 19. The $FOM$ computed for the ECIS is higher than ICIS1-3, SAFAN [31], SIAFA1 [3], SIAFA3 [3], and SIAFA4 [3]. However, it should be emphasized that the $ER_{C_{out}}$ of ECIS is equal to 0, and errors do not propagate from LSBs to MSBs. This full adder’s computations’ accuracy is higher than the other approximate full adders.
The applicability of ECIS is recommended over the other proposed full adders in applications where the accuracy of computations is of great importance and limited reduction of hardware complexity is acceptable.
Table 19: The $FOM$ results of the SIAFA1-4 [3], SAFAN [31], and the proposed approximate full adders in the third scenario’s RCA structure.
Approximate full adder $FOM$
@@ 5 Conclusion
We tried to take a step forward to improve the power wall problem by applying approximate computing. The ICIS1-3 are proposed to improve energy consumption by applying approximate computing and redefining approximate logic from the exact logic method. The ECIS is also proposed to increase the accuracy of computations while improving the circuit analysis metrics compared to the exact full adders. The implementation algorithms of the proposed full adders are designed by applying the IMPLY method as a stateful logic to enable the implementation of these circuits for IMP in the structure of crossbar arrays. ICIS1-3 improve the energy consumption by 21%-41% over approximate SOA and 73%-75% over the exact full adders. ICIS1-3 improve the computational steps by 14%-74% compared to SOA applying the same number of memristors as previous designs. The ECIS also reduces energy consumption by a maximum of 49% and the number of computational steps by 48% compared to exact full adders. The proposed full adders are compared with SOA by error analysis metrics and image quality criteria in three scenarios and three image processing applications. The ECIS has the highest computation accuracy in all three scenarios. ICIS1-3 have acceptable error evaluation criteria in all three scenarios and image processing applications. Circuit evaluation metrics and accuracy in computations are evaluated side by side by presenting a $FOM$. ICIS3, ICIS2, and ICIS1 have the highest ranks among other approximate full adders concerning the $FOM$, respectively.
@@
Acknowledgements
This research did not recieve any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
@@ Acknowledgements
Data is contained within the article.
Conflicts of Interest
The authors declare no conflicts of interest.
Biography
Seyed Erfan Fatemieh was born in Isfahan, Iran, in 1996. He received his B.Sc. degree in Computer Engineering and M.Sc. degree in Computer Architecture from the University of Isfahan, Isfahan, Iran, in 2018 and 2020. He is currently a Ph.D. candidate in Computer Architecture at the University of Isfahan, Isfahan, Iran. His research interests include In-memory Computing, Digital VLSI, Computer Arithmetic, and Quantum Computing and Reversible Circuits.
Mohammad Reza Reshadinezhad was born in Isfahan, Iran, in 1959. He received his B.S. and M.S. degrees from the Electrical Engineering Department of the University of Wisconsin, Milwaukee, USA, in 1982 and 1985, respectively. He has been in the position of lecturer as faculty of computer engineering at the University of Isfahan since 1991. He also received a Ph.D. Degree in computer architecture from Shahid Beheshti University, Tehran, Iran, in 2012. He is currently an Associate Professor in the Faculty of Computer Engineering at the University of Isfahan. His research interests are Digital Arithmetic, Nanotechnology concerning CNTFET, VLSI Implementation, and Cryptography.

### v2

@@ 4.2 Error analysis simulation and comparison
Two parameters should always be evaluated and analyzed in approximate computing. The first one, the circuit analysis metrics improvement, is analyzed in the last subsection. The computation’s accuracy, as the second main parameter in approximate computing, is analyzed in this subsection. Estimating the error magnitude of an error-tolerant application’s outputs is possible by examining error analysis metrics [44]. Error analysis metrics such as MED and NMED can be applied to analyze the proposed circuits’ accuracy. The smaller the error analysis metrics, the higher the computations’ accuracy. All possible input combinations (65536 different input patterns) are applied to the 8-bit approximate adders of scenarios 1-3, designed based on ICIS1-3, ECIS, SIAFA1-4 [3], and SAFAN [34] to calculate the error analysis metrics. The results of error analysis metrics are shown in Table 15.
@@
The main goal of proposing the ECIS is to prevent the spread of inexact $C_{out}$ from LSBs to the MSBs and increase the accuracy of computations. Error analysis metrics (e.g., MED and NMED) are minimum when the ECIS is applied in the structure of scenarios 1-3, according to the results of Table 15. The accuracy of computations increases when the ECIS full adder is applied in larger arithmetic structures compared to the ICIS1-3, SIAFA1-4 [3], and SAFAN [34], based on the results of Table 15. So, the primary purpose of proposing the ECIS is achieved. The error analysis metrics of ICIS2 and ICIS3 full adders are equal, based on their truth tables, in scenarios 1-3, and the results are reported in Table 15. The ED of these two approximate full adders equals 3, and their computations’ accuracy is lower than ECIS (ED=2). The computations’ accuracy decreased by 9%, 2%, and 1%, respectively, by applying the ICIS2 and ICIS3 instead of SIAFA1 [3] and SIAFA3 [3] in the structures of scenarios 1-3. It should be mentioned that the number of computational cycles has also decreased by 5%, 7%, and 9% by applying the ICIS2 and ICIS3 instead of SIAFA1 [3] and SIAFA3 [3] in these three scenarios, respectively. The computations’ accuracy of ICIS1 is lower than the other proposed approximate full adders in this article, SIAFA1 [3], and SIAFA3 [3] in scenarios 1-3. The accuracy of computations decreased from 4%-10% by applying ICIS1 instead of SIAFA1 [3] and SIAFA3 [3] in scenarios 1-3, but the computational steps improved by 5%-9% in these scenarios. This approximate full adder increases the accuracy of computations in scenarios 1-3 compared to SIAFA2 [3], SIAFA4 [3], and SAFAN [34].
@@ 4.3 Application-level simulation and comparison
The reduction of circuit complexity in approximate computing should be accompanied by an acceptable accuracy reduction so that this computational method can be effectively applied in error-resilient applications. Image processing as an error-resilient application is a complex and data-intensive application widely applied in daily human lives [3]. This application is a suitable candidate for in-memory approximate computing by applying memristors [3, 10]. So, the proposed approximate full adders are applied in the computational structures of scenarios 1-3 and are simulated behaviorally in three different image processing applications of image addition, motion detection, and grayscale filter. PSNR, SSIM, and MSSIM are applied to evaluate the output image quality metrics. The proposed full adders have equal single-bit error analysis metrics, but do not have equal output image quality metrics in image processing applications. The reason is that the input distributions in images differ from those in the error analysis simulations [3].
@@ 4.3.1 Image addition application
According to Table 16, the highest quality of the output images is obtained when ECIS is applied in the computational structures of scenarios 1-3 introduced in subsection 4.1. The lowest quality of the output images (PSNR) corresponded to SIAFA2 [3] and SAFAN [34]. Also, the quality of output images is acceptable when ICIS1-3 are applied in the computational structures of scenarios 1-3. The output image quality is higher than 45 dB when the ECIS is applied in scenario 1. The quality of the output images varied between 43.94 dB and 44.5 dB when ICIS1-3, SIAFA1 [3], and SIAFA3 [3] are applied in the computational structure of scenario 1. The quality of the output images is higher than 39 dB only when ECIS is applied in the second scenario’s computing structure. The quality of the output images is between 38-39 dB when ICIS1-3, SIAFA1 [3], and SIAFA3 [3] are applied in the second scenario’s RCA structure. The outputs of the image addition application have the least image quality when SAFAN [34] and SIAFA2 [3] are applied in the second scenario’s arithmetic architecture (PSNR is less than 38 dB). The quality of output images is higher than of the other circuits by applying ECIS and ICIS3 in the third scenario (PSNR is higher than 33 dB). The outputs are acceptable when ICIS1, ICIS2, SIAFA1 [3], SIAFA3 [3], SIAFA4 [3], and SAFAN [34] are applied in the LSBs of the third scenario’s approximate adder. Suppose the number of approximate full adders (ECIS, ICIS1-3, SIAFA1-4 [3], and SAFAN [34]) increases in the LSBs of an 8-bit RCA to six bits. In that case, the image quality of the outputs generated by these RCAs is less than 30 dB, and the output images are unacceptable.
@@ 4.3.2 Motion Detection Application
Two consecutive frames are placed in the subfigures 10(a) and 10(b). The result of the subtraction of these two consecutive frames, applying an exact 8-bit adder is shown in the subfigure 10(c). Subfigures 10(d)-10(g) are the output images obtained from the subtraction of these two consecutive frames applying the proposed approximate full adders in the RCA of scenario 2. The image quality metrics of the output images in different scenarios are written in Table 17.
@@
According to the results of Table 17, ECIS and ICIS1-3 are among the first six approximate full adders in the application of motion detection in all three scenarios. When the proposed circuits are applied in the computational structures of scenarios 1-3, the quality of the output images is higher than 45 dB, 40 dB, and 34.9 dB, respectively. The image quality metrics of the outputs of motion detection application are not acceptable if the number of approximate full adders in the LSBs of an 8-bit approximate adder structure is changed from 5 to 6 (only the two MSBs are computed applying the exact full adder). The PSNR of output images is higher than 29.5 dB, but the motions could not be detected by applying all four proposed approximate full adders in this structure. The same results are obtained by applying the state-of-the-art [3, 34] in this RCA, and it is impossible to detect motion by applying these approximate full adders.
@@ 4.4 Analysis of the trade-off between Circuit and Error evaluation metrics
The proposed circuits were analyzed only by circuit analysis metrics in subsection 4.1. Different error analysis criteria and image quality metrics were applied to analyze the proposed circuits in subsections 4.2 and 4.3. The reduction of accuracy in computations leads to the decrease in circuit complexity in approximate computing. We compared the proposed circuits by simultaneously considering the circuit and error evaluation criteria in this subsection. The $FOM_{1}$ introduced in this article considers the $\underline{E}$nergy consumption and computational $\underline{D}$elay $\underline{P}$roduct (EDP) as circuit analysis metrics and the accuracy-related metrics of 1-NMED and PSNR according to the third scenario’s results.
This $FOM_{1}$ is defined as
$FOM_{1}=\frac{EDP}{(1-NMED)\times PSNR}$ (34)
The results of $FOM_{1}$ for the proposed circuits, SIAFA1-4 [3], and SAFAN [34] are presented in Table 19. The average PSNR value of three 8-bit image processing applications assessed in the last subsection is considered to calculate the $FOM_{1}$. ICIS3, ICIS2, and ICIS1 made the best trade-off between the circuit complexity and the accuracy of the computations, respectively, according to Table 19. The $FOM_{1}$ computed for the ECIS is higher than ICIS1-3, SAFAN [34], SIAFA1 [3], SIAFA3 [3], and SIAFA4 [3]. However, it should be emphasized that the $ER_{C_{out}}$ of ECIS is equal to 0, and errors do not propagate from LSBs to MSBs. This full adder’s computations’ accuracy is higher than the other approximate full adders.
The applicability of ECIS is recommended over the other proposed full adders in applications where the accuracy of computations is of great importance and a limited reduction of hardware complexity is acceptable.
Table 19: The $FOM_{1}$ results of the SIAFA1-4 [3], SAFAN [34], and the proposed approximate full adders in the third scenario’s RCA structure.
Approximate full adder $FOM_{1}$
@@
Energy saving is of great importance in the design of approximate adders. Examining the trade-off between accuracy, energy saving criteria, and improving computational delay (delay saving) in Error-tolerant applications can provide a suitable quantification between the basic objectives of approximate computing. Accordingly, $FOM_{2}$ has been investigated to examine this trade-off based on what has been proposed in [46]. In this criterion, power saving percentage, delay saving percentage, and $PSNR^{2}$ have been applied to simultaneously analyze the improvement of circuit evaluation criteria and approximate output quality (see (35)). The results of $FOM_{2}$ computation for the third scenarios’ approximate adder are written in Table 20.
$FOM_{2}=Energy~saving\times Delay~saving\times PSNR^{2}$ (35)
Table 20: The $FOM_{2}$ results of the SIAFA1-4 [3], SAFAN [34], and the proposed approximate full adders in the third scenario’s RCA structure.
Approximate full adder $FOM_{2}$
SIAFA1 [3] 183.102
SIAFA2 [3] 101.546
SIAFA3 [3] 180.42
SIAFA4 [3] 174.169
SAFAN [34] 184.944
ICIS1 219.784
ICIS2 242.164
ICIS3 245.104
ECIS 100.483
According to the results of Table 20, it can be concluded that ICIS3, ICIS2, and ICIS1 have made the best trade-off between circuit analysis metrics and accuracy in image processing applications.
@@ 5 Conclusion
We tried to take a step forward to improve the power wall problem by applying approximate computing. The ICIS1-3 are proposed to improve energy consumption by applying approximate computing and redefining approximate logic from the exact logic method. The ECIS is also proposed to increase the accuracy of computations while improving the circuit analysis metrics compared to the exact full adders. The implementation algorithms of the proposed full adders are designed by applying the IMPLY method as a stateful logic to enable the implementation of these circuits for IMP in the structure of crossbar arrays. ICIS1-3 improve the energy consumption by 21%-41% over approximate state-of-the-art and 73%-75% over the exact full adders. ICIS1-3 improve the computational steps by 14%-74% compared to state-of-the-art, applying the same number of memristors as previous designs. The ECIS also reduces energy consumption by a maximum of 49% and the number of computational steps by 48% compared to exact full adders. The proposed full adders are compared with state-of-the-art by error analysis metrics and image quality criteria in three scenarios and three image processing applications. The ECIS has the highest computation accuracy in all three scenarios. ICIS1-3 have acceptable error evaluation criteria in all three scenarios and image processing applications. Circuit evaluation metrics and accuracy in computations are evaluated side by side by presenting two $FOMs$. ICIS3, ICIS2, and ICIS1 have the highest ranks among other approximate full adders concerning the $FOM_{1}$ and $FOM_{2}$, respectively.
Appendix A Supplementary application-level simulation
Appendix A reports the results of the supplementary application-level simulation (image addition and grayscale filter). The purpose of this simulation was to gain more confidence in the applicability of the proposed circuits (ICIS1-3 and ECIS) in the above-mentioned applications. Accordingly, the proposed approximate full adders were applied in the third scenario structure, which has the highest degree of approximation among the others. Ten grayscale images in the image addition application and five RGB images in the grayscale filter application were applied to the approximate adder of the third scenario, consisting of the proposed approximate full adders. The image quality evaluation criteria were recorded separately for each circuit, and the average values obtained from the simulations were calculated for each criterion. The results are reported in Tables A1 (image addition application) and A2 (grayscale filter).
Table A1: Image quality metrics of supplementary application-level simulation: image addition application.
Approximate PSNR SSIM MSSIM
full adder (dB)
Scenario 3: three most significant full adders are exact.
ICIS1 31.8257 0.8941 0.8955
ICIS2 32.8602 0.8881 0.8898
ICIS3 32.986 0.8898 0.8908
ECIS 33.5536 0.9064 0.9051
Table A2: Image quality metrics of supplementary application-level simulation: grayscale filter.
Approximate PSNR SSIM MSSIM
full adder (dB)
Scenario 3: three most significant full adders are exact.
ICIS1 30.8058 0.8262 0.886
ICIS2 34.14 0.8957 0.9337
ICIS3 34.0658 0.8886 0.9328
ECIS 35.7012 0.9031 0.941
Based on the results reported in Tables A1 and A2, it can be concluded that all four proposed circuits exhibit appropriate computational accuracy when applied to the third scenario structure, and the outputs produced in image addition and grayscale filter applications are of suitable quality. Considering Table A1 and comparing it with Table 16, it can be concluded that the outputs obtained from applying approximate full adders ICIS1-3 in the third scenario structure follow the same trends in terms of PSNR in the image addition application. In the grayscale filter application, similar behaviors are observed in the approximate full adders ICIS2 and ICIS3, and their image quality is comparable, as shown in Table 18. This trend also applies to Table A2. ICIS1 and ECIS have the lowest and highest image quality, respectively, according to Tables 18 and A2, and the position of both remains the same according to the simulations performed.
Statements & Declarations
Funding
No funding was received for this research.
Competing interests
The authors have no relevant financial or non-financial interests to disclose.
@@ Appendix A Supplementary application-level simulation
Data sharing is not applicable to this article as no new data were created or analyzed in this study.